OpenStudy (anonymous):

write y=(x-4)(x+2) in standard form

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

y = x^2 + 2x -4x - 8 x^2 + 2x - 8 would you like to do it from there to change it into standard form?

OpenStudy (anonymous):

its a trick question?

OpenStudy (anonymous):

idk is that the standard form?

OpenStudy (anonymous):

yes!!! ok I thought you didn't know what the standard form of a quadratic equation will be so was asking a trick question.

OpenStudy (anonymous):

lol okay i get ya now :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it says to use the standard form to change to vertex form and show your work

OpenStudy (anonymous):

well the vertex form is y = a(x - h)^2 + k and the standard form was y = ax^2 +bx + c so you already found the x and a which is like in this equation a = 1 b = 2 c = -8 so if you want the vertex you need to know this equation -b / 2a

OpenStudy (anonymous):

if you do that you will find the k and then when you plug in the k into the standard form equation you will get the h